If almonds sell at $1.20 per pound and walnuts sell for $0.75 per pound, how many pounds of each must be used to make 45 pounds of a mixture to sell at $1.00 per pound?
Answer by Paul(988)

120x+75(45-x)=100(45)
120x+3375-75x=4500
45x=1125
x=25
45-25=20
Hence. for $1.20 it should be 25 and for $0.75 it should be 20.